OpenStudy (anonymous):

The sum of two consecutive odd integers is 36. Find the integers.

OpenStudy (anonymous):

1521

OpenStudy (anonymous):

15 and 21*

OpenStudy (anonymous):

17 and 19

OpenStudy (anonymous):

consecutive

OpenStudy (anonymous):

x+2+x+4=36. 2x+6=36 2x=30 x=15. 15+2=17 15+4=19 Answers: 17 and 19.